8/11 divided by 4/9 estimate the quotient

5/7 times 10/11 estimate the product

Answer by rfer(16322)

(8/11)/(4/9)
(8/11)*(9/4)=72/44=1 7/11
---------------------------
5/7*10/11=50/77